OpenStudy (anonymous):

For what numbers a and b is the system of equations below satisfied by x=b and y=a? x-4y =a 3x +y= 96 b=___________? a=________________?

OpenStudy (anonymous):

line didnt work on the other problem either direct

hero (hero):

b - 4a = a 3b + a = 96 b = 5a 3(5a) + a = 96 15a + a = 96 16a = 96 a = 96/16 a = 6 b = 30

hero (hero):

The first step is to replace x with b and y with a, then solve.

OpenStudy (anonymous):

that was correct thank you

hero (hero):

Of course it was correct @cwhutche